Ask
your own question, for FREE!
Mathematics
7 Online
Julio is drawing tiles from a bag of tiles like the one shown below. He draws a tile, replaces it, and draws another tile. What is the probability he will draw a red tile on both draws? 1/25 16/225 2/21 1/9
Still Need Help?
Join the QuestionCove community and study together with friends!
@Mason_Jones573
@ganeshie8 @Loser66
@nincompoop
There are five red and a total of fifteen tiles. So the probability of getting a red tile on the first attempt is 5/15 = 1/3 Since Julio replaces it before drawing the next time, so the probabilities do not change, and hence getting a red tile is 1/3. The two events - drawing a tile each time - are independent of each other and hence the probability of getting two tiles is a product of the two individual probabilities, which is\[ 1/3 * 1/3 = 1/9 \]
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
7 hours ago
1 Reply
0 Medals
4 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
2 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal